OpenStudy (anonymous):

say that a and b are thte lengths of the legs of a right triangle and c is the length of the hypotenuse. find c if a = 16 and b = 30 a. c=17 b. c=1,156 c. c=34 d. c=68

OpenStudy (anonymous):

Just use:\[c^2 = a^2 + b^2\]

OpenStudy (anonymous):

\[c = \sqrt{16^2 + 30^2} = \sqrt{1156}\] Can you Solve it further or not??

OpenStudy (anonymous):

idk how.

OpenStudy (anonymous):

\[\sqrt{1156} = \sqrt{34 \times 34}\] Can you proceed further or not??

OpenStudy (anonymous):

34

OpenStudy (anonymous):

34

OpenStudy (anonymous):

c

OpenStudy (anonymous):

Yes dear...